OpenStudy (destinyyyy):

Solve the polynomial equation by factoring and then using the zero-product principle.

OpenStudy (destinyyyy):

x^3 +x^2 = 4x+4

OpenStudy (anonymous):

start by making one side 0 by subtracting 4x and 4 from both sides

OpenStudy (anonymous):

\[x^3+x^2−4x−4=0\] Now group the terms and factor out the gcf from each group \[(x^3+x^2)+(−4x−4)=0\]

OpenStudy (destinyyyy):

Okay

OpenStudy (anonymous):

what's the gcf of x³ and x²?

OpenStudy (destinyyyy):

x

OpenStudy (anonymous):

not just x. They both have x². What do you have inside the parentheses if you factor out x²?

OpenStudy (destinyyyy):

Um x

OpenStudy (anonymous):

Not just x. There are 2 terms originally, so there will still be 2 terms after you factor. let's try it this way